OpenStudy (anonymous):

Help with Algebra Pre-test!!!! MEDAL WILL BE AWARDED! Please please please please help!

OpenStudy (anonymous):

4. –2(m – 30) = –6m (1 point) –15 –13 –8 8 5. 2.9n + 1.7 = 3.5 + 2.3n (1 point) 3 –3 .3 1 For questions 6–7, simplify the expression. 6. –6 – 7(c + 10) (1 point) 64 – 7c –76 – 7c 4 – 13c –16 – 13c 7. 3x + 4(x – 6) – 3(x – 7) (1 point) 4x – 45 4x – 3 10x – 3 10x + 45 For questions 8–10, solve the inequality. 8. p + 4 < –24 (1 point) p < –20 p < 28 p < –28 p < 20 9. d over thirteen ≥ –3 (1 point) d ≥ –16 d ≥ 39 d ≥ 10 d ≥ –39 10. –5p > –30 (1 point) p > 6 p < 6 p > –35 p < 35 11. Which inequality matches the graph? number line (1 point) 2n + 5 ≤ 7 2n – 5 < 7 2n + 5 < 7 2n + 5 ≥ 7 12. Which inequality matches the graph? number line (1 point) 6w + 6 > 18 6w + 6 ≥ 18 6w – 6 > 18 –6w + 6 > 14 For questions 13–14, write and solve an equation. 13. A baseball bat and 8 baseballs cost $163.25. The bat costs $106.45. What is the cost c of one baseball? (1 point) 163.25 – 8c = 106.45; $33.71 8c + 106.45 = 163.25; $7.10 8c + 163.25 = 106.45; $7.10 8c – 106.45 = 163.25; $33.71 14. Jacob weighs 64 pounds. Together, he and his sister weigh eight pounds more than two times the weight of his sister. What is the weight, w, of Jacob’s sister? (1 point) 64 + w = 2w + 8; 56 pounds 64 + w – 8 = 2w; 62 pounds 64 + w = 2w; 64 pounds 64 + w = 2w – 8; 72 pounds 15. Paula sold 35 packages of wrapping paper for this year’s Booster Club fundraiser. This is one less than twice the number of packages that she sold last year. How many packages did Paula sell last year? (1 point) 17 packages 16 packages 20 packages 18 packages For question 16–18, write and solve an inequality. 16. An airline requires carry-on luggage to weigh at most 40 pounds. Your suitcase currently weighs 10 pounds. How many pounds p are available for you to fill your suitcase with other items? (1 point) p – 10 ≤ 40; p ≤ 50 40 – p < 10; p < 30 p + 10 ≥ 40; p ≥ 30 10 + p ≤ 40; p ≤ 30 17. Atlanta, Georgia, receives an average of 27 inches of precipitation per year, and there has been 9 inches so far this year. How many more inches, p, of precipitation can Atlanta get and to stay at or below the average? (1 point) p + 9 > 27; p > 18 9 + p ≥ 27; p ≥ 18 p + 9 < 27; p < 18 9 + p ≤ 27; p ≤ 18 18. Chris earns $7.00 per hour working on the weekends. He needs at least $210.00 for a new cell phone. How many hours, h, does Chris need to work on the weekends to buy a new phone? (1 point) h over seven point zero zero > 210.00; h > 30; 30 hours h over seven point zero zero < 210.00; h < 30; 30 hours 7.00h ≥ 210.00; h ≥ 30; 30 hours 7.00h ≤ 210.00; h ≤ 30; 30 hours

OpenStudy (anonymous):

Dude pretests only count for 1 pt. not worth much at all!!! it doesn't even increase your grade

OpenStudy (anonymous):

Yes but my teacher told me that if i dont make a good grade on this then she is failing me for lack of participation (because of things ive done in the past) And i know nothing about this!! I NEED HELP! AND ITS 12am!!!

OpenStudy (anonymous):

4. –2(m – 30) = –6m -2m+60=-6m put all m in one side and the constant numbers in the other side -2m+6m=-60 4m=-60 divide all by 4 to make m alone m=-15

OpenStudy (anonymous):

2.9n + 1.7 = 3.5 + 2.3n this equation is like in the first qoestion make all variables "n" in the left side then the constant num in the other side ~remember that if u change the place of the number or variables in the other side it will change the sign 2.9n-2.3n=3.5-1.7 0.6n=1.8 divide all by 0.6 n=3

OpenStudy (anonymous):

did u understand>?

OpenStudy (anonymous):

yes.. im just so tired :/ sorry.. its 12am and i have to get the rest of this done by tomorrow.. im exhausted and cant think clearly :/

OpenStudy (anonymous):

But thank u for ur help(:

OpenStudy (anonymous):

are u already answer the other numbers?